OpenStudy (anonymous):

find the perpendicular line of 3q-2p=-8;(1,2)

OpenStudy (anonymous):

first, find the slope of the existing line: 1) -2p=-8-3q p=3/2 q +4 so, the slope m =3/2

OpenStudy (anonymous):

perpendicular line has the slope m1=-1/m or m1=-2/3

OpenStudy (anonymous):

3) now write the equation of perpendicular line @ point (1,2): y-2=-2/3 (x-1) y=-2/3 x +2/3 +2 y=-2/3 x +8/3

OpenStudy (anonymous):

thanks so very much